OpenStudy (anonymous):

Ellen wants to purchase a book that is regularly priced at $18. The book is discounted 15%. She also needs to pay a 6% sales tax on the discounted price. What is the total amount Ellen will pay for the book?

OpenStudy (anonymous):

@aahmed123 i wont ignore u

OpenStudy (anonymous):

thank uuu T.T

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

i got 14.38

OpenStudy (anonymous):

but can u help me tho...@aahmed123

OpenStudy (anonymous):

k, in turn can u help me on a question in earth science?

OpenStudy (anonymous):

ok deal

OpenStudy (anonymous):

...somones replying woop

OpenStudy (anonymous):

can you show me the answer choices?

OpenStudy (anonymous):

there are no choices...

OpenStudy (anonymous):

i think its $16.22...

OpenStudy (anonymous):

but idk

OpenStudy (anonymous):

T.T then meh cant be sure, meh got 16.39 T.T

OpenStudy (anonymous):

=.= all of our answers are different, were hopeless =.=

OpenStudy (anonymous):

oh i can tag somone here!

OpenStudy (anonymous):

@Pinkiexoxo WE NEED U DESPERATLY

OpenStudy (anonymous):

Let me see!! :D

OpenStudy (anonymous):

16.22

OpenStudy (anonymous):

knew it....

OpenStudy (anonymous):

im glad i got a human calculator at mai side...

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

Find out how much a 15% discount would subtract from the cost of an $18 book. Subtract how much that discount would be from $18. That's your new price of the book. Once you find the new price, find out how much 6% of that price is, then add that to your discounted price.

OpenStudy (anonymous):

np u should be thanking pinkie though THANK U PINKIE!

OpenStudy (anonymous):

it was incorrect....

OpenStudy (anonymous):

=.=.......

OpenStudy (anonymous):

You shouldn't have clicked it yet! I was working it out! :P

OpenStudy (anonymous):

Super sorry about that!

OpenStudy (anonymous):

:bangs head on table:

OpenStudy (anonymous):

T.T whyyyyyyyyy?

OpenStudy (anonymous):

:nods:

OpenStudy (anonymous):

thank god for the back button..;

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

Ok im going to science tag me Aysha

OpenStudy (anonymous):

hontoni gomenasai, let pinkie work it out forst and lets see

OpenStudy (anonymous):

kk

OpenStudy (jdoe0001):

well... to clarify it a bit

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ 15}}\%\ of\ 18 \implies \cfrac{{\color{brown}{ 15}}}{100}\cdot 18\implies {\color{brown}{ 15}}\div 100\cdot 18\implies \square \\ \quad \\ \\ \quad \\ 18-\square \\ \quad \\ {\color{brown}{ 6}}\%\ of\ 18-\square \implies \cfrac{{\color{brown}{ 6}}}{100}\cdot 18-\square \implies {\color{brown}{ 6}}\div 100\cdot (18-\square) \\ \quad \\ \textit{total amount}\implies 18-[(18-\square)-({\color{brown}{ 6}}\div 100\cdot (18-\square) ) ]\)

OpenStudy (anonymous):

2.70,....15.30,...

OpenStudy (anonymous):

i got 0.92......

OpenStudy (anonymous):

dont ask....

OpenStudy (anonymous):

i failed it im hopeless

OpenStudy (jdoe0001):

well.... 2.7 is correct, that's 15% of 18 so you'd end up with 15.30 as the discounted price then get 6% of 15.30, wich is 0.918 so 18 - (2.70) - (0.918)

OpenStudy (jdoe0001):

or 0.92 is fine for a rounded 0.918

OpenStudy (anonymous):

oh thank god i got it right

OpenStudy (anonymous):

14.38

OpenStudy (jdoe0001):

hmmm hold the mayo

OpenStudy (anonymous):

im holding it its halfway to the bread

OpenStudy (jdoe0001):

hmmm the 6% is tax so that should be ADDED, not subtracted since Ellen is getting the book for 15.30 but paying out of her pocket 6% of that

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

wait what?

OpenStudy (jdoe0001):

so... it should be \(\bf \textit{total amount}\implies 18-[(18-\square)+({\color{brown}{ 6}}\div 100\cdot (18-\square) ) ]\)

OpenStudy (anonymous):

th box is 2.70 right?

OpenStudy (jdoe0001):

15.30 is the discounted price of the book now, 6% of 15.30 is 0.918 or about 92 cents so, she has to pay, the discounted price PLUS the tax or 15.30 + 0.918 <--- her total price

OpenStudy (anonymous):

16.22 if i round the 0.918

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

so thats the answer? just to make clear

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

ok thank you

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

do you mind if i tag you for another question? because i got it wrong and i dont understand why

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

sure, just post anew

OpenStudy (anonymous):

ok thank you